# Pau S. Pujolas

### v1

@@ 1 Introduction
“When a country (USA) is losing many billions of dollars on trade with virtually every country it does business with, trade wars are good, and easy to win," U.S. President Donald J. Trump on Twitter, Mar 2, 2018.
Debate over the importance of international trade imbalances remains a perennial topic in political debates and the media, with large trade deficits frequently cited as justification for protectionist policy measures. [^1] Delpeuch, Fize, and Martin (2024) observe this connection is more than just rhetoric, finding that bilateral imbalances are associated with an increase in protectionist measures. This connection was particularly evident in the recent U.S.-China trade war, in which the trade imbalance between the United States and China was cited as one of the primary motivations for the initial round of tariff escalations. [^2] The text reads “China has consistently taken advantage of the American economy […] The United States has run a trade in goods deficit with China for years, including a \$375 billion deficit in 2017 alone.” Trump (2018). Despite the centrality of trade imbalances in political discourse and public perception of the trade war, the issue has received comparatively little attention in academic studies. [^3] For example, two of the most cited review studies evaluating the economic impact of the U.S.-China trade war contain no references to or discussion of trade imbalances or deficits (Amiti, Redding, and Weinstein, 2019; Fajgelbaum and Khandelwal, 2022).
In this paper we study how trade imbalances impact the welfare effects of trade policy, which we examine both theoretically and quantitatively. To illustrate the interaction between trade policy and trade imbalances, we revisit the classic result on optimal tariff policy by Johnson (1953) and extend it to a context with trade deficits. We demonstrate that a trade deficit makes foreign demand more inelastic, enabling the home country to better manipulate the terms of trade. This interaction affects the welfare impact of tariffs, resulting in higher optimal, or welfare-maximizing, tariff rates. Through numerical simulations, we show that these findings hold both when bilateral trade imbalances lead to aggregate trade imbalances and when there are only bilateral — but not aggregate — trade imbalances.
@@
We explore the quantitative importance of the interaction between trade imbalances and optimal tariff policy within a multi-region, multi-sector applied general equilibrium model featuring input-output linkages and endogenous bilateral trade deficits.444 Allen, Arkolakis, and Takahashi (2019) develop general conditions for computing counterfactual predictions for the class of general equilibrium gravity models, which similarly allow for trade deficits. They show that counterfactual predictions for these models rely only on demand and supply elasticities. They estimate the impact of a flat 10 percent increases in bilateral iceberg trade frictions between the U.S. and China and find small decreases in real output prices and larger decreases in real expenditures. They note, however, that models that include tariffs do not generally fall within the class of universal gravity models, since tariffs serve as an additional source of domestic income. As discussed in Kehoe, Pujolàs, and Rossbach (2017), these models remain the principal framework used by policymakers to assess the economic impact of changes in international trade policy. Our motivating application is a trade war between the United States and China, inspired by the eruption in trade tensions between the two countries that started when, in January 2018, the United States announced tariff hikes on imports from China. This conflict is notable for our framework because China is the largest exporter to the United States, the United States is the largest trading partner of China, and the trade imbalance between them is the largest across all bilateral country pairs.
@@
We calibrate our model to match observed trade flows and tariffs in the data and compute optimal unilateral and Nash equilibrium tariffs between the United States and China. We find that the United States experiences welfare gains in both the unilateral and Nash equilibrium cases compared to the pre-trade war baseline. This result is driven both by the large bilateral deficit that the United States has with China, and by China’s relatively high baseline tariffs on imports from the United States — both of which reduce the potential loss from retaliation. Compared to a baseline with free trade, however, the United States experiences small welfare losses. We reach this conclusion only when both services and input-output linkages are included in our model. This result highlights the importance of conducting trade policy analysis in a fully-specified framework, as emphasized by Caliendo and Parro (2015). We further evaluate the welfare impact of trade wars with other trading partners, and find that the United States would benefit from starting trade wars — by which we mean moving to the Nash equilibrium tariff rates — with a number of its trading partners. These results might help rationalize the increasingly antagonistic stance of countries with deficits in international trade disputes (Delpeuch, Fize, and Martin, 2024).
@@
The reason relatively little attention has been given to the interaction between trade imbalances and international trade policy is likely because trade imbalances are understood first and foremost as a dynamic consideration that arises from movements in domestic fiscal policy, rather than a phenomenon to be explained by international trade policy (Krugman, 1986). [^5] Dix-Carneiro et al. (2023) and Dix-Carneiro and Traiberman (2023) are notable exceptions that focus on the connection between trade imbalances and labor markets. In support of such a view, Furceri et al. (2018) use a reduced form approach to study the impact of the U.S.-China trade war on a variety of outcomes, including the trade balance, and find that the trade war did little to impact trade imbalances. This result is consistent with our modeling approach, and does not imply that trade imbalances cannot moderate the effect of trade policy.
@@
In dynamic models, aggregate trade imbalances arise endogenously as a means for countries to share risk and smooth consumption across periods. [^6] See Backus, Kehoe, and Kydland (1992) and Backus, Kehoe, and Kydland (1994) for early quantitative models to feature endogenous trade imbalances, and Perri and Fogli (2014), Kehoe, Ruhl, and Steinberg (2018), or Steinberg (2019) for more recent work. These frameworks, however, typically do not distinguish between aggregate and bilateral trade deficits, and imbalances are assumed to balance out over time, which may limit the room for trade policy to interact with transient trade imbalances. For instance, Beshkar and Shourideh (2020) study optimal inter-temporal trade policy in a two-country dynamic Ricardian model, in which trade imbalances arise endogenously and function to smooth consumption over time. They show optimal trade policy is counter-cyclical, with optimal tariffs and capital controls in each period uniquely determined by relative productivities in that period. In their framework, trade imbalances are not generally persistent, do not impact the share of production consumed at home and abroad in a given period, and have no direct impact on optimal trade policy. In contrast, in a static framework, trade imbalances necessarily correspond to imbalances in the shares of production consumed across countries. A dynamic model that features persistence in trade imbalances, similar to the persistence we observe in the data (see Lubik and Sablik, 2017; or Kehoe, Ruhl, and Steinberg, 2018), might lead to trade imbalances impacting consumption shares and thereby create space to derive results similar to those in our static framework.
@@
To our knowledge, ours is the first paper to provide optimal tariffs in a fully-specified applied general equilibrium framework with many sectors and input-output linkages across sectors. This is likely due to the computational complexity of calculating Nash equilibria in higher dimension models. For example, He et al. (2017) compute optimal tariffs between two countries in a two-sector, three-country framework where the third country is a ROW aggregate. In contrast, our model has 18 regions and 22 tradable goods-producing sectors. In such a framework, it is unfeasible to compute optimal tariffs through an exhaustive search, since the strategy space grows polynomially with the number of sectors. For example, if we discretized the solution space into 10 possible tariff rates per sector, a three sector model has 1000 possible strategies for each country, whereas a 22 sector model has $10^{22}$ or 10 Sextillion such strategies. [^7] In our application, we discretize the solution space into 40,000 possible tariff rates for each sector. We employ a genetic algorithm, which is a global optimization heuristic, to find optimal tariff rates for each country in response to the other country’s tariffs. We then iterate these best-response tariff rates until we converge to a Nash equilibrium, where neither country wants to deviate. While optimal tariff rates vary across sectors and specifications, in our application to the U.S.–China trade war, we find the average optimal tariff rate is somewhat stable across specifications and in the range of 7–16 percent. This range is generally consistent with optimal tariff rates found elsewhere in the literature, for example by Caliendo et al. (2023).
@@
After evaluating the impact of a trade war in which tariffs are set to maximize welfare, we use our model to measure the welfare implications of the 2018 tariff escalations between the United States and China observed in the data. We find that both the United States and China experienced welfare losses from their recent trade war. [^8] In Appendix B.1 we quantitatively evaluate the United States setting a uniform tariff rate on its trading partners, a proposed policy during the 2024 presidential election. We find that the United States gains from unilaterally tariffs; however, if countries respond to U.S. tariffs in equal measure, all countries experience welfare losses from the resulting global trade war. These losses are consistent with a large number of papers that have calculated the impact of the 2018 tariffs, such as Amiti, Redding, and Weinstein (2019), Flaaen, Hortaçsu, and Tintelnot (2020), or Carroll and Hur (2023). [^9] We refer readers to Fajgelbaum and Khandelwal (2022) for a broad overview of the trade conflict and a review of the numerous studies of its economic impact, the majority of which find small welfare losses for both the United States and China. We find that U.S. tariff increases during the trade war were negatively correlated with the increases consistent with a move towards Nash equilibrium tariff rates. This provides suggestive evidence in favor of arguments by Fetzer and Schwarz (2021) and Kim and Margalit (2021) who claim that the tariffs were politically, rather than economically, motivated.
2 Global Trade Imbalances
Prior to examining the implications of trade imbalances for trade policy, we document patterns in the evolution of global trade imbalances over time. We report bilateral imbalances in each period as the absolute value of net exports divided by gross trade flows, where all objects are measured for a given bilateral relationship. Specifically we compute,
Figure 1: Bilateral Imbalances as Percentage of Gross Trade Flows
Figure 1 shows the evolution of bilateral imbalances, reported as percentages, from 1827–2020, according to three databases: the TRADHIST database constructed by Fouquin and Hugot (2016), the UN Comtrade database, and the BACI databate described by Gaulier and Zignago (2010); [^11] For the Comtrade data, we use the trade flow as reported by the importer, except where not available. The BACI database implements various techniques meant to reconcile discrepancies in reported trade flows in the underlying Comtrade data. where the latter two series are each accessed through the CEPII Gravity Database (Conte et al., 2022). For each series, we compute the weighted average of equation (1) across all bilateral pairs in a given year, where we weight by destination GDP, and then compute a 5 year moving average. [^12] We linearly interpolate logged GDP to fill in years with missing GDP data for the TRADHIST database.
Although it is important to note that the TRADHIST database is an unbalanced panel, meaning the changing composition of the database can play a role in movements across time, Figure 1 shows that bilateral imbalances generally grew from the 1830s up until World War 1. Following the war, global imbalances fell starkly as the United States and other countries moved towards protectionism, as evidenced through the tariff increases in the Fordney–McCumber and Smoot-Hawlley Tariff Acts in 1922 and 1930, respectively (Hayford and Pasurka, 1991). Global imbalances experienced a large spike during World War 2, after which they returned to their pre-war lows. Following the post-war period, global imbalances have increased and are at levels comparable to historic highs. Although the datasets give us a limited window for examining the implications of recent increases in global trade tensions, the plateau or reversal in global imbalances over the past decade is minimal in the context of broader trends. [^13] The 5-year moving averages remain largely unaffected if we exclude 2020 to avoid the fall in trade that coincided with the global COVID-19 pandemic.
These results are consistent with those of Eugster et al. (2020) and Cuñat and Zymek (2023), who find that bilateral imbalances are large and persistent. These papers examine the extent to which trade imbalances respond to or can be explained by trade policy — a topic largely orthogonal to our focus on how trade imbalances alter the welfare implications of trade policy. For example, Cuñat and Zymek (2023) examine the unobserved iceberg trade costs necessary to rationalize bilateral trade imbalances and evaluate the welfare implications of removing these trade barriers. Unlike tariffs, iceberg trade costs do not generate income and therefore have have markedly different welfare implications. Eugster et al. (2020) argue that changes in bilateral imbalances are best explained by macro factors and are not responsive to changes in trade policy. They focus on explaining changes over 5 year periods between 1995–2015, which Figure 1 shows is a period with steadily increasing trade imbalances and is also a period over which tariffs were generally decreasing.
[…]
Figure 2 shows the evolution of aggregate trade imbalances according to the same three datasets we used to report bilateral imbalances. We again weight imbalances by destination GDP when computing the average value in a given year, before plotting the 5 year moving average. Note that equation (2) can be decomposed into two effects, first is the aggregate trade imbalance as a percentage of total trade flows, and the second is trade as a percentage of GDP. Therefore, this figure captures both that net importers (exporters) are importing (exporting) more, relative to their total trade flows, and that trade is increasing as a fraction of GDP. The plot looks similar if one were to plot the aggregate imbalance as a percentage of total trade, rather than GDP, except with a smaller magnitude increase from 1995 onwards.
Although there are small decreases in both bilateral and aggregate trade imbalances coinciding with recent increases in protectionism, their magnitudes are small and fall short from a full reversal. Taken together, these results establish that bilateral and aggregate imbalances are large, and generally growing over the post-war period. These imbalances, which have been associated with increases in protectionism (Delpeuch et al., 2024), motivate our analysis of the interaction between trade imbalances and the welfare impact of tariffs.
@@ 3 Illustrative Two-Country Model
In this section we develop a simple, two-country Armington (1969) model with CES preferences to highlight the interaction between trade imbalances and optimal tariff policy. Our environment is an extension of the canonical framework of Johnson (1953), in which optimal tariffs are shown to be the inverse of the other country’s domestic absorption and trade elasticity — a result we recover when there are no imbalances.
@@
Country 1’s preferences are given by $U^{1}(c_{1}^{1},c_{2}^{1})$, where $c_{1}^{1}$ is Country 1’s consumption of good 1, and $c_{2}^{1}$ is consumption of good 2. All of good 2 is imported, while consumption of good 1 is domestic production minus exports to Country 2, $c_{1}^{1}=E_{1}-c_{1}^{2}$. There is an exogenous trade deficit $d$, which implies that the trade balance in this economy is given by $pc_{2}^{1}=c_{1}^{2}+d$, where $p$ is the price of good 2 in terms of good 1. The government in Country 1 can impose an import tariff $\tau_{1}$, which will affect the relative price of the foreign good.
The price-taking representative consumer in Country 1, faces a tariff on imports, $\tau_{1}$, and solves the following maximization problem,
@@
$\tau_{1}=\frac{1}{\frac{d\log p}{d\log c_{2}^{1}}}.$ (5)

### v2

@@ 1 Introduction
Debate over the importance of international trade imbalances remains a perennial topic in political discourse and the media, with large trade deficits frequently cited as justification for protectionist policy measures. [^1] Delpeuch, Fize, and Martin (2024) observe this connection is more than just rhetoric, finding that bilateral imbalances are associated with an increase in protectionist measures. This connection was particularly evident in the recent U.S.-China trade war, in which the trade imbalance between the United States and China was cited as one of the primary motivations for the initial round of tariff escalations. [^2] The text reads “China has consistently taken advantage of the American economy […] The United States has run a trade in goods deficit with China for years, including a \$375 billion deficit in 2017 alone.” Trump (2018). Despite the centrality of trade imbalances in political rhetoric and public perception of the trade war, the issue has received comparatively little attention in academic studies. [^3] For example, two of the most cited review studies evaluating the economic impact of the U.S.-China trade war contain no references to or discussion of trade imbalances or deficits (Amiti, Redding, and Weinstein, 2019; Fajgelbaum and Khandelwal, 2022).
In this paper we study how trade imbalances impact the welfare effects of trade policy, which we examine both theoretically and quantitatively. To illustrate the interaction between trade policy and trade imbalances, we revisit the classic result on optimal tariff policy by Johnson (1953) and extend it to a context with trade deficits. We demonstrate that a trade deficit makes foreign demand more inelastic, enabling the home country to better manipulate the terms of trade. This interaction alters the welfare impact of tariffs, resulting in higher optimal, or welfare-maximizing, tariff rates. Through numerical simulations, we show that these findings hold both when bilateral trade imbalances generate aggregate trade imbalances and when there are only bilateral — but not aggregate — trade imbalances.
@@
We explore the quantitative importance of the interaction between trade imbalances and optimal tariff policy within a multi-region, multi-sector applied general equilibrium model featuring input-output linkages and endogenous bilateral trade deficits. [^4] Allen, Arkolakis, and Takahashi (2019) develop general conditions for computing counterfactual predictions for the class of general equilibrium gravity models, which similarly allow for trade deficits. They show that counterfactual predictions for these models rely only on demand and supply elasticities. They estimate the impact of a flat 10 percent increases in bilateral iceberg trade frictions between the U.S. and China and find small decreases in real output prices and larger decreases in real expenditures. They note, however, that models that include tariffs do not generally fall within the class of universal gravity models, since tariffs serve as an additional source of domestic income. Unlike tariffs, iceberg trade costs do not generate income and therefore have markedly different welfare implications. As Kehoe, Pujolàs, and Rossbach (2017) discuss, these models remain the principal framework used by policymakers to assess the economic impact of changes in international trade policy. [^5] Our baseline analysis follows the tradition of Applied or Computable General Equilibrium models built using an Armington structure. In Appendix A we repeat our analysis using the multi-sector Eaton and Kortum (2002) Ricardian framework from Caliendo and Parro (2015) and achieve similar results. Our motivating application is a trade war between the United States and China, inspired by the eruption in trade tensions between the two countries that started when, in January 2018, the United States announced tariff hikes on imports from China. This conflict is notable for our framework because China is the largest exporter to the United States, the United States is the largest trading partner of China, and the trade imbalance between them is the largest across all bilateral country pairs.
@@
We calibrate our model to match observed trade flows and tariffs in the data and compute optimal unilateral and Nash equilibrium tariffs between the United States and China. We find that the United States experiences welfare gains in both the unilateral and Nash equilibrium cases compared to the pre-trade war baseline. This result is driven both by the large bilateral deficit that the United States has with China, and by China’s relatively high baseline tariffs on imports from the United States — both of which reduce the potential loss from retaliation. In particular, when we eliminate bilateral trade imbalances between the United States and China, both countries lose from a trade war relative to existing tariffs. Compared to a baseline with free trade, the United States experiences small welfare losses. This conclusion holds only when both services and input-output linkages are included in our model. This result highlights the importance of conducting trade policy analysis in a fully-specified framework, as emphasized by Caliendo and Parro (2015). We further evaluate the welfare impact of trade wars with other trading partners, and find that the United States would benefit from starting trade wars — by which we mean moving to the Nash equilibrium tariff rates — with a number of its trading partners. These results might help rationalize the increasingly antagonistic stance of countries with deficits in international trade disputes (Delpeuch, Fize, and Martin, 2024).
@@
The reason relatively little attention has been given to the interaction between trade imbalances and international trade policy is likely because trade imbalances are understood first and foremost as a dynamic consideration that arises from movements in domestic fiscal policy, rather than a phenomenon to be explained by international trade policy (Krugman, 1986). [^6] Dix-Carneiro et al. (2023) and Dix-Carneiro and Traiberman (2023) are notable exceptions that focus on the connection between trade imbalances and labor markets. Supporting this view, Furceri et al. (2018) use a reduced form approach to study the impact of the U.S.-China trade war on a variety of outcomes, including the trade balance, and find that the trade war did little to impact trade imbalances. This result is consistent with our modeling approach, and does not imply that trade imbalances cannot alter the effects of trade policy.
@@
In dynamic models, aggregate trade imbalances arise endogenously as a means for countries to share risk and smooth consumption across periods. [^7] See Backus, Kehoe, and Kydland (1992) and Backus, Kehoe, and Kydland (1994) for early quantitative models to feature endogenous trade imbalances, and Perri and Fogli (2014), Kehoe, Ruhl, and Steinberg (2018), or Steinberg (2019) for more recent work. These frameworks, however, typically do not distinguish between aggregate and bilateral trade deficits, and imbalances are assumed to balance out over time, which may limit the room for trade policy to interact with transient trade imbalances. For instance, Beshkar and Shourideh (2020) study optimal inter-temporal trade policy in a two-country dynamic Ricardian model, in which trade imbalances arise endogenously and function to smooth consumption over time. They show that optimal trade policy is counter-cyclical, with optimal tariffs and capital controls in each period uniquely determined by relative productivities in that period. In their framework, trade imbalances are not generally persistent, do not impact the share of production consumed at home and abroad in a given period, and have no direct impact on optimal trade policy. In contrast, in a static framework, trade imbalances necessarily correspond to imbalances in the shares of production consumed across countries. A dynamic model that features persistence in trade imbalances, similar to the persistence we observe in the data (see Lubik and Sablik, 2017; or Kehoe, Ruhl, and Steinberg, 2018), might lead to trade imbalances impacting consumption shares and thereby create space to derive results similar to those in our static framework.
@@
To our knowledge, ours is the first paper to provide optimal tariffs in a fully-specified applied general equilibrium framework with many sectors and input-output linkages across sectors. This is likely due to the computational complexity of calculating Nash equilibria in higher dimension models. For example, He et al. (2017) compute optimal tariffs between two countries in a two-sector, three-country framework where the third country is a ROW aggregate. In contrast, our model has 18 regions and 22 tradable goods-producing sectors. In such a framework, it is unfeasible to compute optimal tariffs through an exhaustive search, since the strategy space grows polynomially with the number of sectors. For example, if we discretize the solution space into 10 possible tariff rates per sector, a three sector model has 1000 possible strategies for each country, whereas a 22 sector model has $10^{22}$ or 10 Sextillion such strategies. [^8] In our application, we discretize the solution space into 40,000 possible tariff rates for each sector. We employ a genetic algorithm, which is a global optimization heuristic, to find optimal tariff rates for each country in response to the other country’s tariffs. We then iterate these best-response tariff rates until we converge to a Nash equilibrium, where neither country wants to deviate. While optimal tariff rates vary across sectors and specifications, in our application to the U.S.–China trade war, we find the average optimal tariff rate is somewhat stable across specifications and in the range of 7–16 percent. This range is generally consistent with optimal tariff rates found elsewhere in the literature, for example by Caliendo et al. (2023).
@@
After evaluating the impact of a trade war in which tariffs are set to maximize welfare, we use our model to measure the welfare implications of the 2018 tariff escalations between the United States and China observed in the data. We find that both the United States and China experienced welfare losses from their recent trade war. [^9] In Appendix B.1 we quantitatively evaluate the United States setting a uniform tariff rate on its trading partners. We find that the United States gains from unilaterally imposing such tariffs; however, if countries respond to U.S. tariffs in equal measure, all countries experience welfare losses from the resulting global trade war. These losses are consistent with a large number of papers that have calculated the impact of the 2018 tariffs, such as Amiti, Redding, and Weinstein (2019), Flaaen, Hortaçsu, and Tintelnot (2020), or Carroll and Hur (2023). [^10] We refer readers to Fajgelbaum and Khandelwal (2022) for a broad overview of the trade conflict and a review of the numerous studies of its economic impact, the majority of which find small welfare losses for both the United States and China. We find that U.S. tariff increases during the trade war were negatively correlated with the increases consistent with a move towards Nash equilibrium tariff rates. This provides suggestive evidence in favor of arguments by Fetzer and Schwarz (2021) and Kim and Margalit (2021) who claim that the tariffs were politically, rather than economically, motivated.
@@ 2 Illustrative Two-Country Model
In this section we develop a simple, two-country Armington (1969) model with CES preferences to highlight the interaction between trade imbalances and optimal tariff policy. Our environment is an extension of the canonical framework of Johnson (1953), in which optimal tariffs are shown to be the inverse of the other country’s domestic absorption and trade elasticity — a result that we recover when there are no imbalances.
@@
Country 1’s preferences are given by $U^{1}(c_{1}^{1},c_{2}^{1})$, where $c_{1}^{1}$ is Country 1’s consumption of good 1, and $c_{2}^{1}$ is consumption of good 2. All of good 2 is imported, while consumption of good 1 is domestic production minus exports to Country 2, $c_{1}^{1}=E_{1}-c_{1}^{2}$. There is an exogenous trade deficit $d$, which implies that the trade balance in this economy is given by $pc_{2}^{1}=c_{1}^{2}+d$, where $p$ is the price of good 2 in terms of good 1.111111It is possible for transfers to harm the beneficiary country in general equilibrium models, an effect known as the transfer paradox (see Leontieff, 1936; Samuelson, 1952; or Gale, 1974; among many others). While $d$ in our model appears similar to the transfer in those results, this effect is not what drives our results. In later sections of the paper, we show that our results on optimal tariff setting hold even when we eliminate aggregate trade imbalances, such that only bilateral imbalances remain — and therefore no country receives a “transfer.” The government in Country 1 can impose an import tariff $\tau_{1}$, which will affect the relative price of the foreign good.
The price-taking representative consumer in Country 1 faces a tariff on imports, $\tau_{1}$, and solves the following maximization problem,
@@
$\tau_{1}=\frac{1}{\left(\frac{d\log c_{2}^{1}}{d\log p}\right)}={\frac{d\log p}{d\log c_{2}^{1}}}.$ (3)
@@
3 Global Trade Imbalances
Prior to examining the implications of trade imbalances for trade policy, we document patterns in the evolution of global trade imbalances over time. We report bilateral imbalances in each period as the absolute value of net exports divided by gross trade flows, where all objects are measured for a given bilateral relationship. Specifically we compute
Figure 1: Bilateral Imbalances as Percentage of Gross Trade Flows
Figure 1 shows the evolution of bilateral imbalances, reported as percentages, from 1965–2020, according to two databases: the UN Comtrade database, and the BACI database described by Gaulier and Zignago (2010); [^14] For the Comtrade data, we use the trade flows as reported by the importer, except where not available. The BACI database implements various techniques meant to reconcile discrepancies in reported trade flows in the underlying Comtrade data. where the series are accessed through the CEPII Gravity Database (Conte et al., 2022). For each series, we compute the weighted average of equation (10) across all bilateral pairs in a given year, where we weight by destination GDP, and then compute a 5 year moving average. Figure 1 shows that bilateral imbalances grew throughout the period. Although the datasets give us a limited window for examining the implications of recent increases in global trade tensions, the plateau or reversal in global imbalances over the past decade is minimal in the context of broader trends. [^15] The 5-year moving averages remain largely unaffected if we exclude 2020 to avoid the fall in trade that coincided with the global COVID-19 pandemic.
Our results are consistent with those of Eugster et al. (2020) and Cuñat and Zymek (2023), who find that bilateral imbalances are large and persistent. These papers examine the extent to which trade imbalances respond to or can be explained by trade policy — a topic largely orthogonal to our focus on how trade imbalances alter the welfare implications of trade policy. For example, Cuñat and Zymek (2023) examine the unobserved iceberg trade costs necessary to rationalize bilateral trade imbalances and evaluate the welfare implications of removing these trade barriers. [^16] In Appendix B.2 we consider the implications of iceberg trade costs and their removal in our context of optimal tariff policy. Eugster et al. (2020) argue that changes in bilateral imbalances are best explained by macro factors and are not responsive to changes in trade policy. They focus on explaining changes over 5 year periods between 1995–2015, which Figure 1 shows is a period with steadily increasing trade imbalances and is also a period over which tariffs were generally decreasing.
[…]
Figure 2 shows the evolution of aggregate trade imbalances according to the same two datasets we used to report bilateral imbalances. We again weight imbalances by destination GDP when computing the average value in a given year, before plotting the 5 year moving average. Note that equation (11) can be decomposed into two effects: the first is the aggregate trade imbalance as a percentage of total trade flows, and the second is total trade flows as a percentage of GDP. Therefore, this figure captures that net importers (exporters) are importing (exporting) more relative to their total trade flows, and that trade is increasing as a fraction of GDP. The plot looks similar if one were to plot the aggregate imbalance as a percentage of total trade, rather than GDP, except with a smaller magnitude increase from 1995 onward.
Although there are small decreases in both bilateral and aggregate trade imbalances coinciding with recent increases in protectionism, their magnitudes are small and fall short of a full reversal. Taken together, these results establish that bilateral and aggregate imbalances are large, and generally growing. These imbalances, which have been associated with increases in protectionism (Delpeuch et al., 2024), motivate our analysis of the interaction between trade imbalances and the welfare impact of tariffs.